# Less is More: Improving LLM Alignment via Preference Data Selection

### v3

@@ 4 Experiments
4.1 Experimental Setup
Preference Datasets.
We evaluate our approach using three established preference datasets: (1) Reddit TL;DR summarization dataset (volske2017tl, ; stiennon2020learning, ) that contains human-written summaries and human-rated results, (2) Anthropic Helpful and Harmless dialogue dataset (HH) (bai2022training, ), and (3) UltraFeedback (cui2023ultrafeedback, ), which comprises quality-scored model responses across diverse prompts from multiple sources. To explore how models react to on-policy data, we leverage two modified versions of the UltraFeedback dataset, Llama-UltraFeedback and Mistral-UltraFeedback (meng2024simpo, ). In the variants, the original chosen and rejected responses are replaced with the highest and lowest scored responses, respectively, sampled from five candidates generated by the corresponding Instruct model. The scores are given by the PairRM (jiang2023llm, ) reward model. Statistics about these datasets are in Appendix A.
Models.
Our experiments are conducted across four model series: Llama-3.2 (llama32, ), Llama-3 (dubey2024llama, ), Mistral-7B-v2 (jiang2023mistral, ), and Qwen-2.5 (yang2024qwen2, ) under Base and Instruct setups. For the Base model (Llama-3.2-3B and Llama-3-8B), we first establish fundamental Instruction-following capabilities through supervised fine-tuning on the RLHFlow/SFT-OpenHermes-2.5-Standard datasets. For the Instruct setup, we directly use them as the start of DPO training. Regarding the external reward model, we adopt the recent Skywork-Reward-Llama-3.1-8B-v0.2 (liu2024skywork, ) that is the best reward model at this scale according to the RewardBench leadboard. As for the implicit reward model, we employ the Llama-3.2-3B Base and its DPO-tuned model (on 2,000 randomly selected samples from the complete dataset) for $\pi_{\mathrm{ref}}$ and $\pi_{\theta}$.
Implementation and Evaluation.
For DPO training, we follow (rafailov2024direct, ) and use a fixed value of $\beta=0.1$, except for TL;DR where $\beta=0.5$. We run each training for two epochs, with a learning rate of $5\times 10^{-7}$, and a $0.1$ warmup ratio. Following (rafailov2024direct, ), we evaluate the models using 400 randomly sampled test sets from the validation/test pools of the TL;DR and HH datasets, separately. For models trained on UltraFeedback, we employ AlpacaEval and AlpacaEval 2.0 (li2023alpacaeval, ) as our evaluation benchmark, which consists of 805 diverse questions. [^2] They use the same set of questions, and differ in their reference response generation: AlpacaEval uses Text-Davinci-003 (ye2023comprehensive, ), whereas AlpacaEval 2.0 employs GPT4-1106-preview As the ground truth oracle is unavailable, we employ GPT-4 as a proxy for human judgment across three distinct settings: summarization, helpful or harmless completion, and single-turn dialogue. We utilized a fixed decoding temperature ($T=0.7$) for all model generation in the experiments. More details are presented in Appendix A.
Baselines.
For the offline data selection setting, we compare our method with three types of methods: (1) Random, a simple yet effective strategy in many domains (e.g., Instruction tuning (xia2024less, )), (2) IFD (li2023quantity, ) (i.e., exponential form of the Point-wise Mutual Information), which measures semantic overlap. We use the difference in IFD scores between chosen and rejected responses for preference data selection. (3) External/Implicit Margin (M-Ex/Im) computes the gap between chosen and rejected responses using either external reward models or implicit DPO rewards. For (2) and (3), we segment the data into P (most positive pairs), Z (close to zero pairs), and N (most negative pairs) subsets according to margin values. Specifically, previous work (wu2024beta, ) posits that "hard" preference pairs (where chosen and rejected samples are highly similar) are more beneficial for training, and we use the IFD-Z to quantify this scheme and call it Low-Gap. For the iterative DPO setting, we compare our approach against the standard online iterative DPO baseline established by (xiong2024iterative, ; dong2024rlhf, ) and run for three rounds, each using 20k prompts sampled from UltraFeedback.
@@ 4.2 Win Rate Comparison with Baselines on Classic Preference Datasets
  • Among all methods, only implicit margin-N consistently identifies toxic samples, emphasizing the value of incorporating DPO implicit reward margin into the BeeS strategy. Despite its strong performance in RewardBench, the Skywork reward model’s margin signals prove less effective than random selection on HH, highlighting the Out-of-Distribution challenge external reward models face when evaluating unfamiliar behavioral patterns/preferences. As for the IFD margin metric, it exhibits notable inconsistency across different datasets, rendering it unreliable for evaluating new datasets where prior preference patterns are unknown.
@@
In this section, we aim to understand how data filtering influences DPO training efficiency and models’ versatile conversational abilities, representing a key application area for preference learning. We use both Llama-3-8B (Base) and (Instruct) models, measuring performance through raw and length-controlled win rates on AlpacaEval 2.0, and results are shown in Table 2, LABEL:fig:loss-margin-8b and 6.
@@
Different training dynamics of ‘P’/‘N’/‘Z’ subset region. The left panel of Figure LABEL:fig:loss-margin-8b shows DPO training loss curves for subsets selected by various strategies. Notably, training on subsets filtered according to the ‘P’ and ‘N’ criteria results in a rapid decrease in loss. In contrast, the loss curves corresponding to the ‘Z’ criterion tend to stabilize at consistently higher plateaus. Notably, ’N’-selected samples, which are often assumed as "difficult-to-learn" (gao2025principled, ), can actually be learned as rapidly as ’P’-selected samples, suggesting that ‘bad’ preferences are also easy to grasp for LLM. The pattern is consistent across different datasets and models (see Appendix C.1 for more results. While this observation might explain proposals that use absolute margin values for selection (muldrew2024active, ), Table 1 reveals that ‘P’ and ‘N’ samples produce opposing effects despite similar training dynamics.
The middle and right panels of Figure LABEL:fig:loss-margin-8b illustrate that data subsets selected by BeeS exhibit both the most rapid decrease in training loss and the fastest increase in the DPO training margin, i.e., current train-batch average implicit margin. These concurrent observations of accelerated optimization help to explain the superior performance achieved by BeeS.
@@
Extension to Iterative DPO. We explore the data efficiency of iterative DPO using prompts from UltraFeedback as in (xiong2024iterative, ). In comparison, 20k prompts are used for on-policy preference pair generation per iteration, and our online version uses BeeS to reserve only 5k samples per iteration. The results are in the left panel of Figure 6.
@@
Furthermore, we compared BeeS data selection with several established DPO variants that modify the original loss function, including SimPO (meng2024simpo, ), $\beta$-DPO (wu2024beta, ), and $\alpha$-DPO (wu2024alpha, ). Our findings indicate that BeeS is unique in its ability to effectively enhance both the win rate and the length-controlled (LC) win rate. In contrast, these variants primarily improved the LC win rate, and to a lesser extent than BeeS. These results underscore the significant potential of data selection and data efficiency considerations to enhance the original DPO training algorithm.
@@ 4.4 Ablation Study
BeeS selected subsets are effective for diverse preference learning algorithms. We examine whether the subset selected by BeeS remains data efficient for DPO-variants like IPO (azar2024general, ), KTO (ethayarajh2024kto, ), and SLiC (zhao2023slic, ). We utilize the BeeS selected 6k-sample subset from Llama-UltraFeedback and the results are presented in the right panel of Figure 7. We observe that the high-margin subset consistently benefit these preference learning algorithms by outperforming full-set training. Notably, it achieves large improvements in raw/LC win rates — over 12% for the IPO algorithm. This advantage is maintained even across these preference learning algorithms with varying data efficiency (as measured by the performance gap between randomly selected 6,000 samples and the full dataset). These findings highlight the significant value of sample filtering for other preference learning. Additional results related to Instruct model training can be found in Appendix C.3.
@@ Appendix A Datasets and Evaluation Details
Post: Post
Summary A: Summary_A
Summary B: Summary_B
FIRST provide a one-sentence comparison of the two summaries, explaining which you prefer and why. SECOND, on a new line, state only "A" or "B" to indicate your choice. Your response should use the format: Comparison: <one-sentence comparison and explanation> Preferred: <"A" or "B">
{spverbatim}
@@
Response A: res_a
Response B: res_b
FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful or harmless. SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful or harmless. Your response should use the format: Comparison: <one-sentence comparison and explanation> More helpful or harmless: <"A" or "B">
@@ A.1 More Implementation Details
The baseline strategies are implemented as follows: first, we eliminate outlier samples with extreme margin values (both positively high and negatively low) for CPPL, Ex, and Im metrics. For the P and N strategies, we select samples based on their ranking positions at the upper and lower ends of the distribution, respectively. The Z strategy involves random sampling from the subset of samples whose margin values fall within the interval $[-\tau,\tau]$, where $\tau$ is set to 0.1 for CPPL and 1.0 for Ex/Im metrics.
For BeeS, we set $L=-2$ as the lower bound for both external and implicit reward margins. The upper bound $U$ is determined dynamically based on two conditions: (1) The number of samples with margin values in the interval $[U,\text{max margin}]$ is less than 30, or (2) The number of samples in $[U,\text{max margin}]$ is less than max $\text{max margin}-U$.
@@ B.1 Singular Margin Distribution
The margin distributions calculated using CPPL margin, External and Implicit DPO reward margins, as illustrated in Figures LABEL:fig:cppl-margin,LABEL:fig:ex-margin,LABEL:fig:im-margin, reveal a notable concentration of sample margins around zero. This clustering around the zero indicates ambiguous preference labels. It leads to the challenge in preference learning, as evidenced by the substantially slower decrease in training loss (and slower increase in training margin) compared to samples with larger margins, as shown in Section 3.2.
@@ B.2 Joint Margin Distribution
To complement the left and middle subfigures in Figure LABEL:fig:visual, we present additional results showing the joint margin distributions of samples on the other datasets in Figure LABEL:fig:visual-extra. Our analysis reveals that external and implicit margins exhibit minimal correlation across all four datasets, while implicit margins calculated by different models maintain a high correlation. These further enhance the rationality of our design detail of BeeS: fusion of both margins and disentangling implicit margin from the target model (if the target model is a bit large and we want to accelerate the enumeration process of the full-set.)
@@ C.1 Train Loss and Margin Curves - 3B
To complement the right subfigure in Figure LABEL:fig:visual, we present additional results showing the progression of training loss and margins throughout the DPO training process. The results are shown in Figure LABEL:fig:loss-margin. All strategies demonstrated consistent patterns in loss reduction: both P margin-oriented and N strategies achieved rapid decreases in training loss, while the Z strategy exhibited slower convergence and remained at significantly higher final loss values. Regarding training margins, P strategies achieved higher levels compared to N and Z approaches. Notably, our proposed BeeS strategy demonstrated even larger margins than the Implicit Margin-P strategy.
@@ C.3 More Results for Ablation Study on the DPO Variants
As a complementary study to the results shown in Figure 7, we conducted experiments using the Llama-3-8B-Instruct model while maintaining all other experimental parameters. The results, presented in Figure LABEL:fig:abl_po_ins, demonstrate that models trained on subsets selected by BeeS achieved significantly higher win rates across most evaluation scenarios.

### v4

@@ 4.2 Win Rate Comparison with Baselines on Classic Preference Datasets
  • Among all methods, only implicit margin-N consistently identifies toxic samples, emphasizing the value of incorporating DPO implicit reward margin into the BeeS strategy. Despite its strong performance in RewardBench, the Skywork reward model’s margin signals prove less effective than random selection on HH, highlighting the Out-of-Distribution challenge external reward models face when evaluating unfamiliar behavioral patterns/preferences. As for the IFD margin metric, it exhibits notable inconsistency across different datasets, rendering it unreliable for evaluating new datasets where prior preference patterns are unknown. In general, preference learning departs from traditional representation learning, which predominantly leverages contrastive samples to improve discriminative capacity (Deng et al., 2023b, a). Preference learning focuses on capturing semantic relationships, and models benefit when the underlying preferences in the data are explicit and well-defined.
@@
(a) DPO training loss and margin of Llama-3.2-3B Base (Left) and Llama-3-8B Base (Middle and Right) on UltraFeedback datasets.
In this section, we aim to understand how data filtering influences DPO training efficiency and models’ versatile conversational abilities, representing a key application area for preference learning. We use both Llama-3-8B (Base) and (Instruct) models, measuring performance through raw and length-controlled win rates on AlpacaEval 2.0, and results are shown in Table 3, Figures˜7(a) and 7.
@@
Different training dynamics of ‘P’/‘N’/‘Z’ subset region. The left panel of Figure 7(a) shows DPO training loss curves for subsets selected by various strategies. Notably, training on subsets filtered according to the ‘P’ and ‘N’ criteria results in a rapid decrease in loss. In contrast, the loss curves corresponding to the ‘Z’ criterion tend to stabilize at consistently higher plateaus. Notably, ’N’-selected samples, which are often assumed as "difficult-to-learn" (Gao et al., 2025), can actually be learned as rapidly as ’P’-selected samples, suggesting that ‘bad’ preferences are also easy to grasp for LLM. The pattern is consistent across different datasets and models (see Appendix C.1 for more results. While this observation might explain proposals that use absolute margin values for selection (Muldrew et al., 2024), Table 2 reveals that ‘P’ and ‘N’ samples produce opposing effects despite similar training dynamics.
The middle and right panels of Figure 7(a) illustrate that data subsets selected by BeeS exhibit both the most rapid decrease in training loss and the fastest increase in the DPO training margin, i.e., current train-batch average implicit margin. These concurrent observations of accelerated optimization help to explain the superior performance achieved by BeeS.
@@
Extension to Iterative DPO. We explore the data efficiency of iterative DPO using prompts from UltraFeedback as in (Xiong et al., 2024). In comparison, 20k prompts are used for on-policy preference pair generation per iteration, and our online version uses BeeS to reserve only 5k samples per iteration. The results are in the left panel of Figure 7.
@@
Furthermore, we compared BeeS data selection with several established DPO variants that modify the original loss function, including SimPO (Meng et al., 2024), $\beta$-DPO (Wu et al., 2024b), and $\alpha$-DPO (Wu et al., 2024a). Our findings indicate that BeeS is unique in its ability to effectively enhance both the win rate and the length-controlled (LC) win rate. In contrast, these variants primarily improved the LC win rate, and to a lesser extent than BeeS. These results underscore the significant potential of data selection and data efficiency considerations to enhance the original DPO training algorithm.
@@ 4.4 Ablation Study
BeeS selected subsets are effective for diverse preference learning algorithms. We examine whether the subset selected by BeeS remains data efficient for DPO-variants like IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), and SLiC (Zhao et al., 2023). We utilize the BeeS selected 6k-sample subset from Llama-UltraFeedback and the results are presented in the right panel of Figure 8. We observe that the high-margin subset consistently benefit these preference learning algorithms by outperforming full-set training. Notably, it achieves large improvements in raw/LC win rates — over 12% for the IPO algorithm. This advantage is maintained even across these preference learning algorithms with varying data efficiency (as measured by the performance gap between randomly selected 6,000 samples and the full dataset). These findings highlight the significant value of sample filtering for other preference learning. Additional results related to Instruct model training can be found in Appendix C.3.
@@ 5 Conclusion
Acknowledgements
This work is supported by the National Natural Science Foundation of China (U24B20180) and the Alibaba Group through Alibaba Research Intern Program. We appreciate the reviewers for their insightful feedback and advice, these constructive criticism and recommendations have been invaluable in helping us improve the quality of this work.
@@ Appendix A Datasets and Evaluation Details
Post:
{Post}
Summary A:
{Summary_A}
Summary B:
{Summary_B}
FIRST provide a one-sentence comparison of the two summaries, explaining which you prefer and why. SECOND, on a new line, state only "A" or "B" to indicate your choice. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
Preferred: <"A" or "B">
@@
Response A:
{res_a}
Response B:
{res_b}
FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful or harmless. SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful or harmless. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
More helpful or harmless: <"A" or "B">
@@ A.1 More Implementation Details
The baseline strategies are implemented as follows: first, we eliminate outlier samples with extreme margin values (both positively high and negatively low) for CPPL, Ex, and Im metrics. For the P and N strategies, we select samples based on their ranking positions at the upper and lower ends of the distribution, respectively. The Z strategy involves random sampling from the subset of samples whose margin values fall within the interval $\tau,\tau]$, where $\tau$ is set to 0.1 for CPPL and 1.0 for Ex/Im metrics.
For BeeS, we set L=2L=-2 as the lower bound for both external and implicit reward margins. The upper bound $U$ is determined dynamically based on two conditions: (1) The number of samples with margin values in the interval $U,\text{max margin}]$ is less than 30, or (2) The number of samples in $U,\text{max margin}]$ is less than $\text{max margin}-U$.
@@ Appendix B Visualization of Margin Distributions
(a) Distribution of implicit reward margins on TL;DR, HH, and UltraFeedback datasets. The reward is calculated using the Llama-3.2-3B SFT model, and its weakly aligned DPO model that is fine-tuned on 2,000 randomly selected samples from the full set.
(b) Distribution of external rewards on TL;DR, HH, and UltraFeedback datasets. The reward is calculated using Skywork-Reward-Llama-3.1-8B-v0.2.
(c) Distribution of conditional perplexity (also named instruction following difficulty) margins on TL;DR, HH, and UltraFeedback datasets. The perplexity is calculated with the Llama-3.2-3B SFT model.
@@ B.1 Singular Margin Distribution
The margin distributions calculated using CPPL margin, External and Implicit DPO reward margins, as illustrated in Figures 12(c),12(b),12(a), reveal a notable concentration of sample margins around zero. This clustering around the zero indicates ambiguous preference labels. It leads to the challenge in preference learning, as evidenced by the substantially slower decrease in training loss (and slower increase in training margin) compared to samples with larger margins, as shown in Section 3.2.
@@ B.2 Joint Margin Distribution
To complement the left and middle subfigures in Figure 4(a), we present additional results showing the joint margin distributions of samples on the other datasets in Figure 12(d). Our analysis reveals that external and implicit margins exhibit minimal correlation across all four datasets, while implicit margins calculated by different models maintain a high correlation. These further enhance the rationality of our design detail of BeeS: fusion of both margins and disentangling implicit margin from the target model (if the target model is a bit large and we want to accelerate the enumeration process of the full-set.)
(d) Subfigure (a)-(d): scatter plot showing the joint distribution of samples across external and implicit reward margin values on four datasets. Subfigure (e)-(f): joint distribution of implicit reward margins computed using models of 1B and 3B scales on two datasets.
@@ C.1 Train Loss and Margin Curves - 3B
To complement the right subfigure in Figure 4(a), we present additional results showing the progression of training loss and margins throughout the DPO training process. The results are shown in Figure 12(e). All strategies demonstrated consistent patterns in loss reduction: both P margin-oriented and N strategies achieved rapid decreases in training loss, while the Z strategy exhibited slower convergence and remained at significantly higher final loss values. Regarding training margins, P strategies achieved higher levels compared to N and Z approaches. Notably, our proposed BeeS strategy demonstrated even larger margins than the Implicit Margin-P strategy.
(e) DPO train loss and margin on TL;DR, HH, and UltraFeedback datasets. The training was implemented with Llama-3.2-3B SFT version on different subsets selected by five strategies.
@@ C.3 More Results for Ablation Study on the DPO Variants
As a complementary study to the results shown in Figure 8, we conducted experiments using the Llama-3-8B-Instruct model while maintaining all other experimental parameters. The results, presented in Figure 12(f), demonstrate that models trained on subsets selected by BeeS achieved significantly higher win rates across most evaluation scenarios.
(f) Ablation study on variants of DPO: win rate comparison on IPO, KTO, and SLiC algorithms. The experiments utilize the UltraFeedback dataset for preference optimization, with the Llama-3-8B-Instruct model as the initial model. Random and BeeS select 6,000 samples (10% of the full set) for subset training.
@@ C.4 Hyperparameters Risks
(g) DPO on Llama-UltraFeedback subsets of varying sizes, selected by BeeS. The training is conducted on Llama-3-8B-Base model.